# Grounded GUI Understanding for Vision Based Spatial Intelligent Agent: Exemplified by Virtual Reality Apps

### v3

@@ 4.4. Dataset Statistics
After data cleaning, we finally construct a dataset consisting of 1,552 images with the size of 960 * 540 from 100 apps, covering 245 community-generated genres on Steam, with 4,470 interactable annotations across 766 semantics categories, highlighting the diversity of our dataset. Annotations are in widely used COCO (Lin et al., 2014) format, which uses bounding boxes to mark the locations of IGEs.
@@ 5.1. Research Questions
  • RQ1 (Performance in industrial-setting): How effective is our proposed framework, Orienter, in IGE detection on industrial-setting VR apps?
@@
    • RQ1-3 (Context-sensitive interactability): How does Orienter perform in terms of analyzing context-sensitive interactability?
  • RQ2 (Ablation Study): How does each component of Orienter contribute to its performance?
  • RQ3 (Usefulness): How effectively can Orienter boost automated testing on VR apps?
@@ 5.2. Baselines
GPT-4o (web, 2024b) & Gemini 1.5 Pro (Reid et al., 2024) (direct prompting): Simply prompt the model with the image and ask it to locate all IGEs with bounding boxes and their semantics in the image to illustrate the end-to-end performance of LMM. Xianyu (Chen et al., 2020) is a UI-to-Code tool developed by Alibaba, leveraging old-fashioned computer vision techniques and OCR. UIED (Xie et al., 2020) combines old-fashioned methods with deep learning models to detect clickable GUI elements in complex GUI images. CenterNet2 (Zhou et al., 2021) is a two-stage object detector that uses class-agnostic one-stage detectors as the proposal network. It estimates object probabilities in the first stage and conditionally classify objects in the second stage. Faster R-CNN (Ren et al., 2017) is a two-stage anchor-based deep learning object detector. It contains a region proposal network to extract regions of interest. Objects within the RoIs are then classified with another neural network. YOLO v8 (Jocher et al., 2023) is a one-stage anchor-free detection model. It detects and classifies objects simultaneously and thus is faster than two-stage detectors.
@@ 5.3.1. Implementation of Orienter
For the implementation of Orienter, we leverage the most representative and popular LMMs, i.e., gpt-4-vision-preview, gpt-4o-2024-08-06, Claude 3.5 Sonnet, Gemini 1.5 Pro. For described IGE detection, we leverage the pretrained model of APE-L_D (Shen et al., 2024), which is trained on ten datasets and demonstrates promising results in visual grounding. We apply post-processing to improve Orienter’s prediction quality by filtering abnormally large bounding boxes (over 90% of the image size), and applying Non-Maximum Suppression (NMS) which retains only the highest-confidence box among overlapping ones with IoU exceeds 0.7 to reduce duplicates.
@@ 5.3.2. Dataset Preparation
We partitioned the images into training/validation/testing datasets in a 6:1:3 ratio using three distinct methods: App split, Genre split, and Context-sensitive split. The App split randomly allocates images based on their corresponding apps. The Genre split further considers the apps’ genres to assess the methods’ performance across different app types. The Context-sensitive split assigns images containing the 41 sampled categories in the context dataset to the test set while randomly distributing the rest into the training and validation sets.
@@ 5.3.3. Experimental Setup
We train baselines on the training dataset and evaluate all methods including Orienter on the test dataset. For Xianyu (Chen et al., 2020), we only use its component detection part to find IGEs. For UIED (Chen et al., 2020), we reimplement and train the CNN classifier on our dataset to adapt it to our problem, and replace the OCR component with PaddleOCR. For training on CenterNet2 (Zhou et al., 2021), Faster R-CNN (Ren et al., 2017), and YOLO v8 (Jocher et al., 2023), we set the batch size to 32 and accordingly adjust the learning rate linearly, and keep other configurations the same as their original releases. We apply early stopping with the patience of 20 epochs to avoid overfitting. For LMM experiments, we run three times and regard the average results as final results.
@@
We implement a semantic matching tool to support open vocabulary category matching. We consider two categories semantically match if the cosine similarity of their embedding vectors obtained from the embedding-3 model released by Zhipu AI (Zhipu AI, 2024) exceeds a preset threshold. The first two authors examine the similarity of some typical categories and set the threshold as 0.85.
For RQ1-1, RQ1-2, and RQ2, predictions are evaluated with our customized COCO API that adapts to our semantic matching tool. Note that during evaluation, instead of simply ignoring categories without ground truth annotations (usually due to splitting the dataset), as the official COCO API (web, 2023a) does, we further check if any predictions fall into those categories. Metrics of categories that meet this condition are set to 0 and included when averaging metrics across categories, providing more comprehensive results. Regarding RQ1-3, the baseline methods are trained on the Semantic dataset using Context-sensitive Split. Predictions on the test set are then evaluated on the Context dataset with the same split. The more IGEs and less non-interactable objects the method detects, the better this method’s performance on context-sensitive interactability understanding. Metrics are initially calculated for each category and then averaged to obtain the final results.
5.3.4. VR App Automated Testing Setup
To answer RQ3, we simulate a simplified test scenario: given a screenshot of the scene in the VR app under test, the testing agent attempts to interact with the GUI elements in the scene in a black-box setting. We compare the performance of the testing agent with and without the guidance of Orienter using the test set of Genre split as input.
@@ 5.3.4. VR App Automated Testing Setup
We simplify the interaction events in the 3D virtual space to interaction points on UI images of VR scenes and ignore the type of interactions. The testing process is modeled as incrementally generating points on UI images over time, treating each point as an interaction made by the testing agent with an interval of one minute. We follow the recent works (Pan et al., 2020; Ghorbani et al., 2023) to set the testing duration to 60 minutes, which is longer than exploring a VR scene usually requires. We perform 5 testing runs and average the metrics to produce comprehensive results for each interaction strategy.
@@
For the non-guided strategy, the testing agent randomly generates interaction points on the whole image. For the guided strategy, in each attempt, the testing agent generates interaction within the bounding boxes predicted by Orienter at a probability $p$, or ignores the prediction’s constraint to randomly explore the whole UI images at the probability $(1-p)$. The probability $p$ decreases gradually as the test progresses. Suppose the current moment is $t$ minutes and the total duration of the test is $T$ minutes, then $p=1-t/T$. This strategy allows the agent to generate effective yet diversified interaction events to cover more IGEs precisely.
@@ 5.4. Evaluation Metrics
Precision, Recall, and F1-Score. A predicted bounding box matches the one with the highest IoU among the ground-truth bounding boxes that (1) lie in the same image, (2) have matching categories, and (3) have IoU exceeding the threshold. For RQ1-1, RQ1-2, and RQ2, we consider a predicted box matching a ground-truth box a True Positive (TP); otherwise, a False Positive (FP); ground-truth boxes that fail to match any predicted box are False Negatives (FN). For RQ1-2, a predicted box is considered a TP if it matches an interactable ground-truth box, and an FP if it matches a non-interactable one; non-interactable and interactable ground-truth boxes that fail to match any predicted box are considered a TN and FN, respectively. We calculate Precision as $TP/(TP+FP)$, Recall as $TP/(TP+FN)$ and F1-Score as $2\times Precision\times Recall/(Precision+Recall)$.
@@
We follow the previous works (Wang, 2022; Wang et al., 2023b) to use IGE Coverage and Effective Interaction Count in RQ3. Note that we can not calculate code coverage metrics in the black-box setting of RQ3.
IGE Coverage. The IGE coverage reflects the effectiveness of interaction strategies to find IGEs over time. An IGE is covered if at least one interaction point falls within the bounding box of that IGE. IGE Coverage is calculated as $n_{covered}/n_{all}$, where $n$ represents the number of IGEs.
Effective Interaction Count. The Effective Interaction Count reflects the efficiency of interaction strategies to make effective interactions by focusing on IGEs only. An interaction point is considered effective if it falls in any of IGEs’ bounding boxes.
@@ 6. Results and Analysis
6.1. RQ1: Performance in industrial-setting

### v4

@@ 4.4. Dataset Statistics
After data cleaning, we construct a dataset consisting of 1,552 images with the size of 960 * 540 from 100 apps, covering 12 official app genres and 245 community-generated app categories on Steam, with 4,470 interactable annotations across 766 semantics categories, highlighting the diversity of our dataset. Annotations are in widely used COCO (Lin et al., 2014) format, which uses bounding boxes to mark the locations of IGEs. Detailed distribution for apps in app genres and annotations in IGE’s semantics categories are shown in Figure 13(a) and Figure 13(b), respectively.
(a) # App in different genres. Genres in different color are used as different set in genre splitting of the dataset introduced in section 5.3.2
(b) # Annotation in the First 30 IGE categories associated with the most annotations. The IGE category ”button” which contributes 555 annotations is ommitted
Figure 13. Dataset distribution
@@ 5.1. Research Questions
  • RQ1 (IGE detection performance in industrial-setting): How effective is Orienter in IGE detection on industrial-setting XR apps?
@@
    • RQ1-3 (Context-sensitive interactability): How does Orienter perform in terms of analyzing context-sensitive interactability, i.e. the same kind of object with different interactability in different contexts?
  • RQ2 (Ablation Study): How does each component of Orienter contributes to its performance?
  • RQ3 (Effectiveness for improving testing): How effectively can Orienter boost automated testing on XR apps?
@@ 5.2. Baselines
GPT (web, 2024c) & Gemini (Reid et al., 2024) & Claude (web, 2024b) (direct prompting): Simply prompt the model with the image and ask it to find the location and semantics of all IGEs. Xianyu (Chen et al., 2020) is a UI-to-Code tool leveraging old-fashioned computer vision techniques and OCR. UIED (Xie et al., 2020) combines old-fashioned methods with deep learning models to detect clickable elements in GUI images. CenterNet2 (Zhou et al., 2021) is a two-stage object detector that first estimates object probabilities images and then conditionally classify objects. Faster R-CNN (Ren et al., 2017) is a two-stage anchor-based deep learning object detector with a region proposal network and a classification network. YOLO v8 (Jocher et al., 2023) is a one-stage anchor-free detection model that detects and classifies objects simultaneously. OmniParser (Wan et al., 2024) combines object detection, OCR, and LMM to parse GUI images into element trees for GUI grounding.
@@ 5.3.1. Implementation of Orienter
For the implementation of Orienter, we leverage the most representative and popular LMMs, i.e., gpt-4-vision-preview, gpt-4o-2024-08-06, Claude 3.5 Sonnet, Gemini 1.5 Pro. For described IGE detection, we leverage the pretrained model of APE-L_D (Shen et al., 2024), which is trained on ten datasets and demonstrates promising results in visual grounding. We apply post-processing to improve Orienter’s prediction quality by filtering abnormally large bounding boxes ( $90\%$ image size), and applying Non-Maximum Suppression (NMS) which retains only the highest-confidence box among overlapping ones with IoU exceeds 0.7 to reduce duplicates.
@@ 5.3.2. Dataset Preparation
Table 1. Statistics of different splitting methods. Ctx. is short for Context.
Splitting Method Set # Covered App # GUI Scene Image # IGEs
- Full 100 1,552 4,470
App Train 60 925 2,459
Valid 10 163 574
Test 30 464 1,437
Genre Train 65 958 2,528
Valid 11 136 610
Test 24 458 1,332
Ctx.-sensitive Train 96 947 2,583
Valid 64 158 404
Test 53 447 1,483
Test
(Ctx. dataset)
53 447 1,239
$\bigcup${App, Genre} Test 46 687 2,034
$\bigcup${Ctx.-sensitive,
App, Genre}
Test 77 936 2,875
We partition the images into training/validation/test sets in a ratio of approximately 6:1:3 using three distinct splitting methods: App splitting, Genre splitting, and Context-sensitive splitting. App splitting randomly allocates images based on their corresponding apps such that images from the same app are placed in the same set. Genre splitting further considers the apps’ genres to assess the methods’ performance across different app types. Context-sensitive splitting assigns images containing the 41 sampled categories in the context dataset to the test set while randomly distributing the remaining images into the training and validation sets.
The statistics of each splitting method are illustrated in Table  1. The usage of the datasets and splitting methods will be discussed in section 5.3.3 and section 5.3.4.
@@ 5.3.3. Experimental Setup
We train baselines on the training set and evaluate all methods including Orienter on the test set. For Xianyu (Chen et al., 2020), we only use its component detection part to find IGEs. For UIED (Chen et al., 2020), we train the CNN classifier on our dataset to adapt it to our problem. For training on CenterNet2 (Zhou et al., 2021), Faster R-CNN (Ren et al., 2017), and YOLO v8 (Jocher et al., 2023), we set the batch size to 32 and accordingly adjust the learning rate, apply early stopping with the patience of 20 epoch. Other configurations are the same as their original releases. For LMM experiments, we run three times and regard the average results as the final results.
@@
In terms of evaluation, we implement a semantic matching tool to support open vocabulary category matching. We consider two categories semantically match if the cosine similarity of their embedding vectors obtained from the embedding-3 model released by Zhipu AI (Zhipu AI, 2024) exceeds a preset threshold. The first two authors examine the similarity of some typical categories and set the threshold to 0.85. We accrodingly modify the official COCO API (web, 2023a) to support semantic matching.
For RQ1-1, RQ1-2, and RQ2, predictions are evaluated with the modified COCO API. We leverage App splitting and Genre splitting to evaluate each method in RQ1-1 and RQ1-2, then use the union of both splitting’s test set to assess the contribution of each components of Orienter in RQ2.
Regarding RQ1-3, the baseline methods are trained on the Semantic dataset using Context-sensitive Splitting. Predictions on the test set are then evaluated on the Context dataset with the same splitting. The more IGEs and less non-interactable objects the method detects, the better this method’s performance on context-sensitive interactability understanding.
5.3.4. XR App Automated Testing Setup
@@
To answer RQ3, we simulate a simplified test scenario: given a screenshot of the scene in the XR app under test (AUT), the testing agent attempts to interact with the GUI elements in the scene in a black-box setting. We compare the performance of the testing agent with and without the guidance of Orienter using the union of test sets of the 3 splitting as input, regarding the covered 77 apps as AUTs. The testing process is modeled as incrementally generating points on UI images over time, treating each point as an interaction made by the testing agent with an interval of one minute. We follow the recent works (Pan et al., 2020; Ghorbani et al., 2023) to set the testing duration to 60 minutes, which is longer than exploring a XR scene usually requires. We perform 30 testing runs and average the metrics for each interaction strategy.
@@
For the non-guided (i.e. random) strategy, the testing agent randomly generates interaction points on the whole image. For the guided strategy, in each attempt, the testing agent generates interaction within the bounding boxes predicted by Orienter at a probability $p$, or randomly explore the whole UI images at the probability (1p)(1-p). The probability $p$ decreases gradually as the test progresses. Suppose the current moment is $t$ minutes and the total duration of the test is $T$ minutes, then p=1t/Tp=1-t/T. This strategy allows the agent to generate effective yet diversified interaction events to cover more IGEs precisely.
@@ 5.4. Evaluation Metrics
Precision, Recall, and F1-Score. A predicted bounding box matches the one with the highest IoU among the ground-truth bounding boxes that (1) have matching categories, and (2) have IoU exceeding the threshold. For RQ1-1, RQ1-2, and RQ2, we consider a predicted box matching a ground-truth box a True Positive (TP); otherwise, a False Positive (FP); ground-truth boxes that fail to match any predicted box are False Negatives (FN). For RQ1-2, a predicted box is considered a TP if it matches an interactable ground-truth box, and an FP if it matches a non-interactable one; non-interactable and interactable ground-truth boxes that fail to match any predicted box are considered as TNs and FNs, respectively. We calculate $Precision=TP/(TP+FP)$, $Recall=TP/(TP+FN)$ and $F1=2\times Precision\times Recall/(Precision+Recall)$.
@@
We follow previous works (Wang, 2022; Wang et al., 2023b) to use IGE Coverage and Effective Interaction Count in RQ3.
IGE Coverage. The IGE coverage reflects the effectiveness of strategies to find IGEs over time. An IGE is covered if at least one interaction point falls in the bounding box of that IGE. IGE Coverage is calculated as $n_{covered}/n_{all}$, where $n$ represents the number of IGEs.
Effective Interaction Count. The Effective Interaction Count reflects the efficiency of strategies to make effective interactions by focusing on IGEs only. An interaction point is considered effective if it falls in any of IGEs’ bounding boxes.
@@ 6. Results and Analysis
6.1. RQ1: IGE detection performance in industrial-setting
Methods Interactability Semantics
App Splitting Genre Splitting App Splitting Genre Splitting
Faster-
RCNN
mAP% 18.05 13.84 7.32 1.47 0.07 17.90 12.88 6.96 1.56 0.13 0.21 0.15 0.03 0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01
P% 33.41 32.29 23.04 11.08 2.15 33.42 29.55 20.10 8.06 1.14 0.91 0.83 0.38 0.16 0.06 0.07 0.06 ¡0.01 ¡0.01 ¡0.01
R% 32.00 25.00 18.00 8.00 1.00 29.00 22.00 15.00 6.00 1.00 4.00 4.00 3.00 3.00 4.00 2.00 2.00 ¡0.01 ¡0.01 ¡0.01
F1% 32.69 28.18 20.21 9.29 1.36 31.05 25.22 17.18 6.88 1.06 1.48 1.37 0.68 0.30 0.11 0.13 0.12 ¡0.01 ¡0.01 ¡0.01
Center-
Net2
mAP% 17.79 14.98 11.73 5.91 0.93 15.63 12.53 9.07 4.23 0.45 0.20 0.18 0.15 0.05 ¡0.01 0.22 0.19 0.05 0.02 ¡0.01
P% 32.53 28.52 26.85 18.26 4.83 28.67 25.81 27.71 16.46 4.51 0.68 0.65 0.57 0.27 0.05 0.62 0.59 0.27 0.19 0.04
R% 32.00 29.00 23.00 15.00 3.00 30.00 27.00 18.00 12.00 3.00 4.00 4.00 3.00 3.00 2.00 5.00 5.00 3.00 3.00 3.00
Yolo v8 mAP% 15.30 13.12 9.85 6.16 1.51 11.54 10.02 7.59 3.86 0.46 0.02 0.02 0.01 ¡0.01 ¡0.01 0.03 0.02 0.02 0.01 0.01
P% 26.58 25.38 20.63 16.65 7.31 23.92 21.38 24.22 14.55 2.68 0.21 0.20 0.19 0.05 0.01 0.29 0.28 0.27 0.25 0.21
R% 26.00 23.00 20.00 14.00 3.00 23.00 22.00 15.00 11.00 3.00 4.00 4.00 4.00 1.00 3.00 3.00 3.00 3.00 3.00 3.00
OmniParser mAP% 0.13 0.13 0.13 0.05 ¡0.01 0.29 0.28 0.21 0.08 0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01
P% 2.48 2.44 2.22 2.37 0.29 6.19 5.90 2.93 1.76 0.96 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01
R% 2.00 2.00 2.00 1.00 ¡0.01 3.00 3.00 3.00 2.00 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01
F1% 2.21 2.20 2.10 1.41 ¡0.01 4.04 3.98 2.96 1.87 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01
Gemini 1.5
Pro
mAP% ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 0.11 ¡0.01 ¡0.01 ¡0.01 ¡0.01 0.12 ¡0.01 ¡0.01 ¡0.01 ¡0.01
P% 0.21 ¡0.01 ¡0.01 ¡0.01 ¡0.01 0.43 0.43 ¡0.01 ¡0.01 ¡0.01 0.25 ¡0.01 ¡0.01 ¡0.01 ¡0.01 0.23 ¡0.01 ¡0.01 ¡0.01 ¡0.01
R% ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 4.00 ¡0.01 ¡0.01 ¡0.01 ¡0.01 50.00 ¡0.01 ¡0.01 ¡0.01 ¡0.01
F1% ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 0.47 ¡0.01 ¡0.01 ¡0.01 ¡0.01 0.46 ¡0.01 ¡0.01 ¡0.01 ¡0.01
gpt-4o-
2024-08-06
mAP% 0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01
P% 0.78 ¡0.01 ¡0.01 ¡0.01 ¡0.01 0.75 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01
R% ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01
F1% ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01
Orienter
(gpt-4-vision-
preview)
mAP% 25.53 24.32 23.03 19.64 4.61 27.07 25.22 23.87 19.80 4.54 14.93 14.66 13.96 12.17 3.67 15.96 15.57 14.62 12.87 3.75
P% 34.79 33.52 32.32 29.31 14.92 38.42 36.70 34.97 31.40 16.79 16.06 16.05 15.17 15.27 6.53 17.11 16.85 15.78 15.93 6.14
R% 40.00 39.00 38.00 34.00 12.00 40.00 38.00 37.00 32.00 11.00 50.00 50.00 50.00 33.00 20.00 50.00 50.00 50.00 33.00 20.00
Orienter
(gpt-4o-
2024-08-06)
mAP% 26.44 25.26 23.98 20.35 4.48 26.11 24.50 23.51 19.42 4.38 18.00 17.65 16.91 14.28 5.75 19.45 18.97 18.07 15.40 5.68
P% 34.63 34.89 35.43 31.64 18.57 37.91 36.68 36.71 33.24 12.91 19.32 18.97 18.28 15.37 6.87 20.35 19.81 19.01 16.19 7.36
R% 42.00 39.00 36.00 34.00 11.00 41.00 39.00 37.00 33.00 14.00 50.00 50.00 50.00 50.00 42.00 50.00 50.00 50.00 50.00 25.00
Orienter
(Claude 3.5
Sonnet)
mAP% 23.99 22.90 21.54 18.46 3.90 25.50 23.96 22.21 18.49 3.84 16.14 15.79 15.17 12.73 5.01 19.01 18.51 17.78 15.00 6.67
P% 31.73 37.52 36.79 33.56 16.30 37.97 40.15 39.22 35.86 14.51 16.63 16.28 15.83 13.35 7.64 19.68 19.11 18.53 15.75 9.87
R% 40.00 32.00 31.00 29.00 11.00 37.00 33.00 32.00 29.00 13.00 50.00 50.00 50.00 50.00 20.00 50.00 50.00 50.00 50.00 20.00
Orienter
(Gemini 1.5
Pro)
mAP% 32.71 31.27 30.51 26.33 5.88 35.03 32.59 30.82 25.08 5.64 20.69 20.35 20.19 17.77 7.16 22.12 21.44 21.17 18.45 7.41
P% 46.42 45.30 44.94 41.64 21.46 49.45 47.52 46.42 41.94 19.02 22.64 22.29 22.16 19.39 9.66 23.72 22.86 22.64 19.58 8.35
R% 42.00 41.00 40.00 37.00 14.00 43.00 42.00 41.00 37.00 15.00 50.00 50.00 50.00 50.00 25.00 50.00 50.00 50.00 50.00 50.00
    The highest metrics are colored with gray. All metrics of UIED and Xianyu are lower than 0.01% and omitted.